# VocabTailor: Dynamic Vocabulary Selection for Downstream Tasks in Small Language Models

### v1

@@ 3.3 The Hybrid Vocabulary Selection Strategy
3.4 Fine-grained Construction of the Static Task Vocabulary
Our static vocabulary construction algorithm addresses the challenge of isolating truly task-essential tokens through a three-stage filtering process:
Input-Aware Filtering
We first exclude all input tokens from the candidate vocabulary, isolating tokens the model must generate without input cues (e.g., function keywords in code generation, discourse markers in summarization). This step directly implements task-driven locality by identifying the irreducible core $\mathcal{T}$ that cannot be derived from the input context.
Language-Specific Filtering
To suppress noise in mixed-language scenarios (e.g., code datasets with multilingual comments), we apply Unicode block analysis to retain only tokens belonging to the target language family. This heuristic-based approach effectively handles cross-lingual interference while preserving task-relevant vocabulary.
Tolerance Filtering
When additional vocabulary reduction is required, we formulate vocabulary selection similar to a pruning problem. We define the impact metric $df(v)$ as the document frequency—the fraction of instances where token $v$ appears in the ground truth. Tokens are sorted by ascending $df(v)$ and iteratively removed until the cumulative document frequency of pruned tokens reaches a user-defined tolerance threshold $\tau$. This threshold bounds the worst-case performance drop, where $\tau=0.01$ ensures that at most 1% of profiling samples lose a critical token.
4 Experiments
4.1 Settings
Tasks and datasets
@@ 4.1 Settings
Evaluation metrics
We include sacreBLEU (Post 2018), METEOR (Banerjee and Lavie 2005), and COMET (Rei et al. 2020) for machine translation. For summarization, we use Rouge-1, Rouge-2, and Rouge-L scores (Lin 2004). Pass@1 (Chen et al. 2021) is used for code completion. We use F1 score for information extraction, and accuracy for math problem solving. These metrics are standard in the field and provide a robust measure of model performance across the target task.
Models
For machine translation, we use Qwen3-1.7B (Yang et al. 2025). For summarization and information extraction, we employ Llama 3.2 1B (Dubey et al. 2024). For summarization, the base model is fine-tuned for a better base performance. For code completion, we choose deepseek-coder-1.3b-base (Guo et al. 2024), and for math problem solving, we apply rho-math-1b-interpreter-v0.1 (Lin et al. 2024).
Baselines and Other Settings
We compare our method (VocabTailor) with the original model (Original) and static vocabulary pruning (VP). Static vocabulary pruning follows the common routines of corpus-based filtering. For a fair comparison, VP and VocabTailor use the same profiling corpra. We set the tolerance threshold $\tau=0.01$ on all tasks for VocabTailor. As a hybrid dynamic-static framework, the vocabulary size of VocabTailor model varies in each single inference. Here we report the average vocabulary size for each downstream task. For all models, we set the temperature to 0 to avoid randomness.
4.2 Results
Task Model Vocabulary Metric
@@ 4.2 Results
VP 64,117 (42.28%) 14.840/11.842/81.191
VocabTailor (Ours) 18,874 + [40] (12.47%) 15.393/12.692/81.440
@@
VP 65,518 (43.21%) 22.585/46.735/74.933
VocabTailor (Ours) 24,185 + [14] (15.96%) 21.127/46.970/75.493
Summarization Original 128,000 0.364/0.147/0.292
VP 59,613 (90.99%) 0.364/0.147/0.292
VocabTailor (Ours) 36,332 + [15] (55.48%) 0.364/0.147/0.292
Code Completion Original 32,000 54.097%
VP 24,888 (77.78%) 8.124%
VocabTailor (Ours) 3,521 + [58] (11.18%) 53.865%
Information Extraction Original 128,000 38.400
VP 49,106 (38.36%) 2.580
VocabTailor (Ours) [105] (0.08%) 62.730
Math Problem Solving Original 32,000 88.40
VP 10,300 (32.19%) 87.80
VocabTailor (Ours) 5,135 + [14] (16.09%) 88.40
Machine Translation
[…]
Summarization
For summarization, all three models yield identical ROUGE scores, indicating no difference in summary quality. However, VocabTailor achieves this with just 55% of the original vocabulary, 35% fewer tokens than VP. These results demonstrate that aggressive vocabulary reduction is possible without compromising output quality when apply a hybrid dynamic-static vocabulary selection.
Code Completion
On the SAFIM benchmark with deepseek-coder-1.3b-base, VP leads to a dramatic drop in performance: Pass@1 falls from 54.097% to 8.124%, despite a relatively modest vocabulary reduction (down to 77.78% of the original size). This highlights a key limitation of corpus-based pruning for code: essential elements such as variable names and identifiers may be discarded if they appear infrequently in the profiling corpus. In contrast, VocabTailor retains only 11.18% of the vocabulary and yet achieves a high Pass@1 of 53.865%. These findings underscore the robustness of our input- and task-specific vocabulary selection strategy for generation-intensive tasks such as code synthesis.
Information Extraction
For information extraction tasks, we evaluate on the SQuAD dataset using LLaMA 3.2 1B. VocabTailor achieves the most striking result: with just 0.08% of the original vocabulary retained (and no static tokens at all), it attains an F1 score of 62.73, outperforming both the Original (38.4) and VP (2.58) by large margins. This result stems from the nature of extractive tasks, where the output vocabulary is typically a subset of the input. Because VocabTailor preserves input tokens dynamically, it retains all the necessary vocabulary for accurate extraction. The poor performance of VP suggests that static, corpus-based pruning is poorly suited for tasks where input-output overlap is high, whereas VocabTailor is especially effective.
Math Word Problem Solving
For symbolic reasoning, we evaluate on a math problem-solving task. The unpruned model achieves a score of 88.40, which is fully preserved under VocabTailor, even though it reduces the vocabulary to just 16% of the original. VP, while also reducing vocabulary size to 32%, causes a slight performance drop to 87.80. This shows that even in tasks requiring high-precision symbolic handling, our approach remains effective.
4.3 Ablation Study
Impact of Dynamic vs. Static Vocabulary Components
We first evaluate the impact of the dynamic and static vocabulary components, both individually and in combination. As shown in Table 2, using only the dynamic part that contains tokens profiled from the specific input examples, results in a significant performance drop (Pass@1 of 36.064%). This demonstrates that input tokens alone lack the broader coverage needed for robust code completion. Using only the static part (task-specific tokens) achieves a Pass@1 of 52.3%. However, this still underperforms the full static-dynamic configuration, which achieves the best result at 53.865% with nearly the same vocabulary size. This indicates that while the static tokens carry most of the task-relevant capacity, including the dynamic tokens adds critical input-specific nuances, and their combination is essential for optimal performance.
We also compare our static-only approach with VP: VP retains 78% of the original vocabulary—more than our approach—it results in a drastically lower Pass@1 of 8.124%. This stark contrast underscores a key insight: while both VP and our static-only setup are static, VP’s direct modification of the tokenizer and embeddings damages input representations, leading to severe degradation. In contrast, our method retains the full tokenizer and embedding, thereby preserving representational integrity and maintaining high performance even with significantly fewer tokens.
Model Vocabulary Pass@1
Original 100% 54.097%
VP 77.80% 8.124%
Dynamic + Static ($\tau=0.01$) 11.18% 53.865%
Dynamic only 0.81% 36.064%
Static only ($\tau=0.01$) 11.00% 52.300%
Impact of Input-aware and Language-specific Filtering
As discussed earlier, VocabTailor calibrates the static vocabulary through three-stage filtering. As shown in Table 3, starting from the unfiltered static vocabulary (78% of the original), applying input-aware filtering (IA) reduces the size to 53% with virtually no performance loss (54.074% vs. 54.062%). Adding language-specific filtering (LS) further reduces the vocabulary size to 46%, while performance slightly improves to 54.085%. This improvement likely stems from the removal of noisy or irrelevant tokens from multilingual corpora, allowing the model to focus on task-relevant representations. These results demonstrate that IA and LS can significantly compress the vocabulary without degrading accuracy, validating the effectiveness of our static token selection process.
Model Vocabulary Pass@1
Original 100.00% 54.097%
Dynamic + Unfiltered static 77.78% 54.074%
Dynamic + IA 52.85% 54.062%
Dynamic + IA + LS 45.61% 54.085%
Impact of Tolerance Filtering
Lastly, we analyze the effect of tolerance filtering, which allows for further reduction of rarely activated tokens based on cumulative document frequency. In Table 4, we vary the tolerance threshold ($\tau$) to observe the trade-off between vocabulary size and accuracy. At $\tau=0$, we preserve all profiled tokens, achieving a Pass@1 of 54.085%. As tolerance increases, more tokens are filtered out, reducing the vocabulary to as low as 2.5% of the original size ($\tau=0.10$), with a gradual decline in performance. Importantly, even with $\tau=0.01$, the vocabulary shrinks to 11.18% with only 1.8% drop in Pass@1, suggesting that our method is robust to aggressive reduction. This highlights the flexibility of tolerance as a tuning knob to balance compression vs. accuracy.
Model Vocabulary Pass@1
$\tau=0$ 45.61% 54.085%
$\tau=0.01$ 11.18% 53.865%
$\tau=0.02$ 7.28% 53.714%
$\tau=0.10$ 2.50% 52.277%
5 Conclusion
6 Limitations & Future Work
The proposed framework presented in this paper is only explored on language downstream tasks of SLMs. This method may be extended and applied to VLMs, ALMs, and MLLMs for memory-efficient inference on image/video understanding and audio generation tasks. Custom kernel implementations and hardware-specific optimizations can be integrated into the framework for further inference optimization.

### v2

@@ Appendix D Profiling Strategy Algorithm
Appendix E Ablation Study in Details
E.1 Impact of Dynamic vs. Static Vocabulary Components
We first evaluate the impact of the dynamic and static vocabulary components, both individually and in combination. As shown in Table 7, using only the dynamic part that contains tokens profiled from the specific input examples results in a significant performance drop (Pass@1 of 36.06%). This demonstrates that input tokens alone lack the broader coverage needed for robust code completion. Using only the static part (task-specific tokens) achieves a Pass@1 of 52.30%. However, this still underperforms the full static-dynamic configuration, which achieves the best result at 53.87% with nearly the same vocabulary size. This indicates that while the static tokens carry most of the task-relevant capacity, including the dynamic tokens adds critical input-specific nuances, and their combination is essential for optimal performance.
Model Vocabulary Pass@1
Original 100% 54.10%
VP 77.80% 8.12%
Dynamic + Static ($\tau=0.01$) 11.18% 53.87%
Dynamic only 0.81% 36.06%
Static only ($\tau=0.01$) 11.00% 52.30%
We also compare our static-only approach with VP: VP retains 78% of the original vocabulary—more than our approach—it results in a drastically lower Pass@1 of 8.12%. This stark contrast underscores a key insight: while both VP and our static-only setup are static, VP’s direct modification of the tokenizer and embeddings damages input representations, leading to severe degradation. In contrast, our method retains the full tokenizer and embedding, thereby preserving representational integrity and maintaining high performance even with significantly fewer tokens.
E.2 Impact of Input-aware and Language-specific Filtering
As discussed earlier, VocabTailor calibrates the static vocabulary through three-stage filtering. In Table 8, starting from the unfiltered static vocabulary (78% of the Original), applying input-aware filtering (IA) reduces the size to 53% with virtually no performance loss (54.07% vs. 54.06%). Adding language-specific filtering (LS) further reduces the vocabulary size to 46%, while performance slightly improves to 54.09%. This improvement likely stems from the removal of noisy or irrelevant tokens from multilingual corpora, allowing the model to focus on task-relevant representations. These results demonstrate that IA and LS can significantly compress the vocabulary without degrading accuracy, validating the effectiveness of our static token selection process.
Model Vocabulary Pass@1
Original 100.00% 54.10%
Dynamic + Unfiltered Static 77.78% 54.07%
Dynamic + IA 52.85% 54.06%
Dynamic + IA + LS 45.61% 54.09%
E.3 Impact of Tolerance Filtering
Lastly, we analyze the effect of tolerance filtering, which allows for further reduction of rarely activated tokens based on cumulative document frequency. In Table 9, we vary the tolerance threshold ($\tau$) to observe the trade-off between vocabulary size and accuracy. At $\tau=0$, we preserve all profiled tokens, achieving a Pass@1 of 54.09%. As tolerance increases, more tokens are filtered out, reducing the vocabulary to as low as 2.5% of the original size ($\tau=0.10$), with a gradual decline in performance. Importantly, even with $\tau=0.01$, the vocabulary shrinks to 11% with only 1.8% drop in Pass@1, suggesting that our method is robust to aggressive reduction. This highlights the flexibility of tolerance as a tuning knob to balance compression vs. accuracy.
  Model   Vocabulary   Pass@1
  $\tau=0$   45.61%   54.09%
  $\tau=0.01$   11.18%   53.87%
  $\tau=0.02$   7.28%   53.71%
  $\tau=0.10$   2.50%   52.28%
Appendix F Data Processing and Evaluation Details
For machine translation, we follow the best practice of zero-shot machine translation by zhang_prompting_2023, using a simple English template.
English-to-Chinese translation template:
@@ Appendix F Data Processing and Evaluation Details
English-to-Italian translation template:
@@
For summarization, the model is finetuned and evaluated using prompt shown below.
Summarization prompt template:
@@
For the code completion task, we use the evaluation script of SAFIM. For information extraction, we use the LM-Eval-Harness framework and set the task to ‘squad_completion‘. For math word problem solving, we use the Math-Eval-Harness framework and set the task to ‘mawps‘.
Appendix G Fine-tuning Config for Summarization Baseline Model
For summarization, all the methods are based on a finetuned Llama 3.2 3B model. We set learning rate=2e-5, batch size = 128. The model is fine-tuned using XSUM train split set for 1 epoch.